# TOPol: Capturing and Explaining Multidimensional Semantic Polarity Fields and Vectors

### v1

@@ Introduction
Sentiment Polarity scoring has long occupied a central position in Natural Language Processing (NLP), treated as a synonym of Semantic Polarity since its inception, addressed within the scope of Sentiment Analysis, conceptualized as a scalar indicating the extent to which a text encodes a positive or negative perspective about an object (Mäntylä, Graziotin, and Kuutila 2018). Recent works dealing with Polarity Classification, still framed in the lens of sentiment, however seek to move beyond affective components for capturing dimensions such as sarcasm (Prasanna, Shaila, and Vadivel 2023). While this sentiment-centric framing of polarity has proven powerful, it overlooks the multidimensional nature of semantic orientation.
@@
In this section, we describe the proposed TOPol framework, as illustrated in Figure 1. The pipeline reconstructs multidimensional semantic polarity fields and interprets the underlying polarity vectors that encode topic-specific semantic shifts contained in the discourse. TOPol begins by generating dense vector representations for each document in the input corpus using a general-purpose transformer model. Next, it applies UMAP (McInnes, Healy, and Melville 2020) on these embeddings, yielding a non-linearly reduced semantic manifold in which document proximity reflects topical similarity. TOPol then partitions the manifold using the Leiden (Traag, Waltman, and van Eck 2019) community detection algorithm, grouping semantically related documents in latent topics. Each topic cluster, or community, is further subdivided into two (or more) context-specific subsets, denoted $A$ and $B$, according to the HoTL-defined contextual boundary. For each cluster, TOPol then computes a polarity vector for each Leiden community by calculating the vector displacement between the centroid of documents in regime $A$ and the centroid of documents in regime $B$. These vectors are then anchored to a shared origin, forming a vector field in a normed vector space, the so-called semantic polarity field for the corpus. Finally, to interpret the polarity dimensions contained in each polarity vector, potentially several per vector, we apply a tLLM-based contrastive explainability procedure. Specifically, for each polarity vector, we provide the tLLM with the document subsets nearest to centroids $A$ and $B$ and prompt it to generate contrastive summaries describing the dominant semantic distinctions between regimes $A$ and $B$, along with their estimated coverage and exemplar text captions.

### v2

@@ Introduction
Sentiment Polarity scoring has long occupied a central position in Natural Language Processing (NLP), treated as a synonym of Semantic Polarity since its inception, addressed within the scope of Sentiment Analysis, conceptualized as a scalar indicating the extent to which a text encodes a positive or negative perspective about an object (Mäntylä et al. 2018). Recent works dealing with Polarity Classification, still framed in the lens of sentiment, however seek to move beyond affective components for capturing dimensions such as sarcasm (Prasanna et al. 2023). While this sentiment-centric framing of polarity has proven powerful, it overlooks the multidimensional nature of semantic orientation.
@@
In this section, we describe the proposed TOPol framework, as illustrated in Figure 1. The pipeline reconstructs multidimensional semantic polarity fields and interprets the underlying polarity vectors that encode topic-specific semantic shifts contained in the discourse. TOPol begins by generating dense vector representations for each document in the input corpus using a general-purpose transformer model. Next, it applies UMAP (McInnes et al. 2020) on these embeddings, yielding a non-linearly reduced semantic manifold in which document proximity reflects topical similarity. TOPol then partitions the manifold using the Leiden (Traag et al. 2019) community detection algorithm, grouping semantically related documents in latent topics. Each topic cluster, or community, is further subdivided into two (or more) context-specific subsets, denoted $A$ and $B$, according to the HoTL-defined contextual boundary. For each cluster, TOPol then computes a polarity vector for each Leiden community by calculating the vector displacement between the centroid of documents in regime $A$ and the centroid of documents in regime $B$. These vectors are then anchored to a shared origin, forming a vector field in a normed vector space, the so-called semantic polarity field for the corpus. Finally, to interpret the polarity dimensions contained in each polarity vector, potentially several per vector, we apply a tLLM-based contrastive explainability procedure. Specifically, for each polarity vector, we provide the tLLM with the document subsets nearest to centroids $A$ and $B$ and prompt it to generate contrastive summaries describing the dominant semantic distinctions between regimes $A$ and $B$, along with their estimated coverage and exemplar text captions.